OpenStudy (anonymous):

how many times does the graph of the function below touch or cross the x-axis? 3x^2+8x+5

OpenStudy (anonymous):

do you know how to do it

hero (hero):

How many times do you think it crosses the x-axis?

OpenStudy (anonymous):

i have no idea

hero (hero):

Use the discriminant test b^2 - 4ac to figure it out. If the discriminant is negative, then it crosses zero times If the discriminant is zero, then it crosses one time. If the discriminant is positive, then it crosses two times

OpenStudy (anonymous):

i got -4

OpenStudy (anonymous):

os my final answer is zero

hero (hero):

a = 3 b = 8 c = 5 b^2 - 4ac = (8)^2 - 4(3)(5) = 64 - 60 = 4 That's what I get

OpenStudy (anonymous):

ok so it zero right i know the answer is 4 but it tells me how many times i cross so the answer is zero

hero (hero):

Re-read what I wrote @romanortiz65

hero (hero):

@DebbieG, allow @romanortiz65 to figure this out on his own.

hero (hero):

@romanortiz65, I'm waiting for you to respond again.

OpenStudy (debbieg):

What is the connection between the discriminant and the touching/crossing behavior? Remember - it's a quadratic, so the shape of the graph is a parabola. If discr>0, there are 2 DISTINCT real roots - what does that mean in terms of it's behavior crossing/touching the axis? If discr=0, there is ONE REPEATED real root - what does that mean in terms of it's behavior crossing/touching the axis? if discr<0, there are NO REAL roots (only complex roots) - what does that mean in terms of it's behavior crossing/touching the axis?

OpenStudy (anonymous):

i got it 2

OpenStudy (debbieg):

@Hero I don't give answers. Was just trying to help him think about it.

hero (hero):

but I was already helping him

OpenStudy (anonymous):

gut no fight aight its cool i got thanks